Find the area of the polygon with the vertices P(-5,4), Q(2,4) and S(1,-1)
Express your answer in the simplest radical form
Please note : Student is only permitted to use formula Area= 1/2 bh
Cannot use this formula Area = 1/2[x1(y2 - y3) + x2(y3 - y1) + x3(y1 - y2)]

Answer by greenestamps(13200)



P and Q have the same y coordinate 4, so PQ is a horizontal segment whose length is the difference between the x coordinates of P and Q: (2-(-5))=2+5=7. Use it as the base.

The height is then the difference between 4 and the y coordinate of S: 4-(-1)=4+1=5.

So the area using 1/2 base times height is (1/2)(7)(5).